# Designing across domains with declarative thinking: Insights from the 96-Eyes ptychographic imager project

### v1

@@ 6.2 Adopts to new biological problems with Math rigor
Continuing the working example, drug designers and microbiologists would point out the need to monitor cell apoptosis under influences of small molecules through the fluorescent-tagged transport protein, which necessities the need of quantitative phase measurements from the time-lapse image frames. Ideally, for monolayer adherent cell lines on an optically flat substrate, the quantitative phase image should exhibit bright-colored cell structures against the black-colored uniform background in grayscale. Ptychographic imaging experts would then raise a concern over heavy signal attenuation of quantitative phase information in the imaging hardware. After all, near-darkfield image is not available and non-negotiable in the 96-Eye project.
@@ A.2 Additional R&D directions uncovered by the problem statement in 5GL form
Domain needs Solutions offered by domain experts Consumable Optics Computer vision (CV) Convex optimization (CVX) Hardware (HW) acceleration Life sciences Biocompatibility of the microwell surface; Tissue culture (TC-)treated surface. Phase-retrieval from intensity signals; Weak-phase assumption Recover adherent cell morphology; Edge-preserving TV regularization in phase channel Quick access to frames from time lapse experiments; Image reconstruction with previous frames as algorithm initial guess. Thermal/Noise / Fan vibration constraints of benchtop instruments; offload memory-intensive pipelines to FPGAs. Consumable Out of focus due to 96-well plate curvature and tip-tilt variations[62]; Ptychographic computational re-focusing of phase channel. Strong phase image background due to uneven TC-treated coating; foreground-background separation prior in the phase channel. Well plate edge effect due to uneven liquid evaporation per well; in situ wavevector calibration via differentiable optics formulation. Optics FPM wavevector calibration error due to liquid meniscus variations; sidewall coating to achieve $90^{\circ}$ liquid contact angle. Errors of the oblique illumination angles due to liquid meniscus; meniscus curvature compensating wavevector calibration. Reconstruction failure due to low-precision optics; computational lens aberration recovery; CV Strong phase image background due to uneven TC-treated coating;[4] develop new TC-treatment assays, use optical quality substrate. TV-regularization prior is compute-intensive in spatial domain; Approximation of the $\nabla$ operator with circular boundary condition Acceleration of advanced, non-linear image denoising priors; HW-accelerated non-local means filter baked onto the GPU silicon CVX Simulating image blur with 2D convolution is compute-intensive; derive the matrix-free operations with Fourier optics. Poor lens aberration estimation due to lack of darkfield images; Zernike wavefront decomposition to improve quality. Migration of CPU-accelerated code to GPUs; decoupling algorithm CAD and CAM with 4GL, e.g. Halide[12].

### v2

@@ 6.2 Adopts to new biological problems with Math rigor
Continuing the working example, drug designers and microbiologists would point out the need to monitor cell apoptosis under influences of small molecules through the fluorescent-tagged transport protein[19], which necessities the need of quantitative phase measurements from the time-lapse image frames. Ideally, for monolayer adherent cell lines on an optically flat substrate, the quantitative phase image should exhibit bright-colored cell structures against the black-colored uniform background in grayscale. Ptychographic imaging experts would then raise a concern over heavy signal attenuation of quantitative phase information in the imaging hardware[7, 20, 21]. After all, near-darkfield image is not available and non-negotiable in the 96-Eye project due to the tall chimneys of the 96-well microwell plates.
@@ A.2 Additional R&D directions uncovered by the problem statement in 5GL form
Domain needs Solutions offered by domain experts
Consumable Optics Computer vision (CV) Convex optimization (CVX) Hardware (HW) acceleration
Life sciences Biocompatibility of the microwell surface; Tissue culture (TC-)treated surface. Phase-retrieval from intensity signals; Weak-phase assumption Recover adherent cell morphology; Edge-preserving TV regularization in phase channel Quick access to frames from time lapse experiments; Image reconstruction with previous frames as algorithm initial guess. Thermal/Noise / Fan vibration constraints of benchtop instruments; offload memory-intensive pipelines to FPGAs.
Consumable Out of focus due to 96-well plate curvature and tip-tilt variations[65]; Ptychographic computational re-focusing of phase channel. Strong phase image background due to uneven TC-treated coating; foreground-background separation prior in the phase channel. Well plate edge effect due to uneven liquid evaporation per well; in situ wavevector calibration via differentiable optics formulation.
Optics FPM wavevector calibration error due to liquid meniscus variations; sidewall coating to achieve $90^{\circ}$ liquid contact angle. Errors of the oblique illumination angles due to liquid meniscus; meniscus curvature compensating wavevector calibration. Reconstruction failure due to low-precision optics; computational lens aberration recovery;
CV Strong phase image background due to uneven TC-treated coating;[4] develop new TC-treatment assays, use optical quality substrate. TV-regularization prior is compute-intensive in spatial domain; Approximation of the $\nabla$ operator with circular boundary condition Acceleration of advanced, non-linear image denoising priors; HW-accelerated non-local means filter baked onto the GPU silicon
CVX Simulating image blur with 2D convolution is compute-intensive; derive the matrix-free operations with Fourier optics. Poor lens aberration estimation due to lack of darkfield images; Zernike wavefront decomposition to improve quality. Migration of CPU-accelerated code to GPUs; decoupling algorithm CAD and CAM with 4GL, e.g. Halide[12].